OpenStudy (anonymous):

Can anyone help me with this? p=920 Length = 100 P= 2L + 2W How do i solve this??

OpenStudy (anonymous):

p=2l+2w 920=2(100)+2w 920=200+2w Now, solve for w. Do you know how? :o)

OpenStudy (anonymous):

No, thats actually what i was trieng to figure out. :/

OpenStudy (anonymous):

Okay, well you're trying to get the w alone on one side of the equation, so you're going to work backwards and do opposite operations. For example, if it's +2 you're going to -2. This will cancel it out. And remember whatever you do to one side you MUST do to the other. I'll show you the first step and you try and do the second. :o) 920=200+2w -200 -200 See how it's a positive 200? That explains the -200. Always take from the side of the equation with the variable (or w in this case). 720=2w See how that cancels the 200 on the right side? Now, notice the 2w is multiplying. So, use the opposite of multiplication to find the value of w. Try it :o)

OpenStudy (anonymous):

So it would be 720=2w -200 -200 ?? o.o im so confused >-<

OpenStudy (anonymous):

No, the 200 is gone now, it canceled out. You have to get w alone, but the 2 is still on the right side of the equation. So, to get w alone, since 2 is multiplying by w, we are going to divide both side by 2. This will cancel the 2 on the right side and leave us with just w. 720=2w /2 /2 360=w Does that make sense?